### v1

# Channel‑Selected Stratified Nested Cross‑Validation for Clinically Relevant EEG‑Based Parkinson’s Disease Detection

Abstract
The early detection of Parkinson’s disease remains a critical challenge in clinical neuroscience, with electroencephalography offering a non‑invasive and scalable pathway toward population‑level screening. While machine learning has shown promise in this domain, many reported results suffer from methodological flaws, most notably patient‑level data leakage, inflating performance estimates and limiting clinical translation. To address these modeling pitfalls, we propose a unified evaluation framework grounded in nested cross‑validation and incorporating three complementary safeguards: (i) patient‑level stratification to eliminate subject overlap and ensure unbiased generalization, (ii) multi‑layered windowing to harmonize heterogeneous EEG recordings while preserving temporal dynamics, and (iii) inner‑loop channel selection to enable principled feature reduction without information leakage. Applied across three independent datasets with a heterogeneous number of channels, a convolutional neural network trained under this framework achieved 80.6% accuracy and demonstrated state‑of‑the‑art performance under held‑out population block testing, comparable to other methods in the literature. This performance underscores the necessity of nested cross‑validation as a safeguard against bias and as a principled means of selecting the most relevant information for patient‑level decisions, providing a reproducible foundation that can extend to other biomedical signal analysis domains.
I Introduction
Parkinson’s disease (PD) is the second most common neurodegenerative disorder, affecting over 10 million people worldwide and imposing a growing socioeconomic burden [1]. Its progressive course—marked by motor dysfunction, cognitive decline, and reduced quality of life—makes early detection and intervention a critical priority  [2, 3]. Yet current clinical diagnosis is largely based on overt motor symptoms, which typically emerge only after substantial dopaminergic neuron loss [2, 3]. This delay highlights the importance of identifying prodromal biomarkers, such as REM sleep behavior disorder, olfactory impairment, and subtle motor irregularities, which often precede clinical diagnosis by years [2]. Detecting early PD could enable interventions when therapies are effective [2].
Advances in biomedical signal acquisition have created new opportunities for non‑invasive monitoring of these prodromal and progressive symptoms [4]. Among available modalities, electroencephalography (EEG) is particularly promising: it is portable, relatively inexpensive, and offers millisecond‑level temporal resolution of neural dynamics [5]. Compared to imaging techniques such as MRI or PET, which provide high spatial resolution but are costly and less accessible, EEG enables scalable, repeated assessments that are well‑suited for longitudinal monitoring [5, 6]. This combination of accessibility and temporal precision positions EEG as a strong candidate for early PD biomarker discovery.
Figure 1: The nested cross‑validation architecture addresses the challenge of high‑dimensional EEG data by starting with all channels and systematically reducing to the most informative. The inner loop selects channels contributing most to inference, while the outer loop validates this choice without bias. This principled “many‑to‑best” reduction is dataset‑, channel‑, and model‑agnostic, yielding results that are more efficient, generalizable, and interpretable.
In parallel, machine learning (ML) and deep learning (DL) methods show strong potential for extracting subtle biomarkers from high‑dimensional biomedical data [7]. A wide range of neural architectures have been explored for biosignal analysis: recurrent neural networks (RNNs) and long short‑term memory (LSTM) models capture temporal dependencies in sequential data [8]; transformer‑based models leverage self‑attention to model long‑range interactions across time and frequency; and graph neural networks (GNNs) exploit the spatial topology of EEG electrode layouts to learn structured representations [9]. Despite these advances, convolutional neural networks (CNNs) remain a principled starting point for EEG analysis. When applied to short‑time Fourier transform (STFT) spectrograms, CNNs exploit partial shift‑invariance that complements the local stationarity of the STFT, enabling robust feature extraction without handcrafted preprocessing [10, 11, 12]. This synergy has consistently delivered state‑of‑the‑art performance across diverse signal domains, including PD detection, where CNN‑based or hybrid pipelines frequently report accuracies exceeding 95% [13]. Moreover, CNN architectures are naturally extensible to multimodal fusion: EEG inputs can be paired with bio-signal sensor data through parallel branches converging in shared layers, allowing integrated biomarker discovery without redesigning the core model [14].
Many reported successes in EEG‑based PD detection are undermined by methodological flaws, most notably data leakage and dataset heterogeneity. Leakage arises in two primary forms: temporal leakage, where overlapping windows from the same recording are split across training and test sets, and subject leakage, where data from the same patient appears in both folds. In both cases, models risk memorizing idiosyncratic traits rather than learning disease‑relevant features, inflating performance and eroding clinical validity [15]. Surveys of PD classification studies further highlight that EEG research often employs ambiguous patient‑level stratification, if it is applied at all [16]. Population block testing offers a partial safeguard, yet its scope is limited, as models may still overfit to the unique characteristics of a single cohort. These concerns are compounded by the heterogeneity of public EEG datasets: channel counts range from high‑density 64‑channel caps to sparse consumer devices, and acquisition hardware differs in sampling rate, electrode placement, and preprocessing pipelines. Methods tuned to one configuration often fail to transfer across others, creating a significant reproducibility barrier. Collectively, these challenges underscore the need for architectures that prevent leakage, generalize across populations, and robustly reduce high‑dimensional signals to the most informative subset, independent of dataset or recording setup.
To address these challenges, we propose a nested cross-validation framework for robust EEG analysis (Figure 1). The outer loop enforces patient-level stratification to eliminate leakage, while the inner loop selects the most informative channels, reducing dimensionality without sacrificing clinically relevant signals. A multi-layered windowing strategy integrates recordings of varying length, sampling rate, and hardware, enabling consistent evaluation across heterogeneous datasets. By design, the framework is dataset, channel, and model-agnostic, positioning it as a blueprint for reproducible EEG-based biomarker discovery. Together, these components yield generalizable performance estimates and clinically meaningful early PD detection results.
Our contributions are threefold:
  1. 1.
    A model‑agnostic nested cross‑validation framework with stratification, windowing, and inner‑loop channel selection to prevent leakage, harmonize datasets, and reduce dimensionality.
  2. 2.
    Empirical evidence that patient‑level leakage and narrow cohort blocking inflate EEG‑based PD detection performance and undermine clinical validity.
  3. 3.
    Demonstration that learned representations align with canonical EEG frequency bands, yielding interpretable models that achieve 0.806 accuracy across datasets.
Looking forward, this framework provides a foundation for accelerating clinical translation: it enables reproducible biomarker validation across heterogeneous cohorts, supports integration with multimodal data streams such as EMG and wearable sensors, and lays the groundwork for scalable digital health pipelines capable of monitoring neurodegenerative disease progression in real-world settings.
II Related Work and Methodological Pitfalls
A growing body of research has applied ML techniques to the early detection of PD using EEG, EMG, and other biosignal modalities. Several studies have reported high classification accuracies, often exceeding 85%, suggesting strong potential for clinical translation [17, 18]. However, closer inspection reveals that many of these works rely on evaluation protocols that do not stratify data at the patient level. Instead, temporal windows from the same subject are distributed across both training and test sets, creating a form of data leakage. This allows models to exploit subject‑specific idiosyncrasies rather than learning generalizable disease markers, thereby inflating performance estimates [15, 19, 20, 21].
Beyond this central issue, several additional pitfalls recur across the literature. The small sample sizes typical of PD datasets magnify overfitting risks, as even a few misallocated subjects can distort reported outcomes [20]. Reported performance is also often reduced to accuracy, which obscures clinically relevant trade‑offs; complementary metrics such as sensitivity, specificity, ROC‑AUC, and calibration curves provide a more reliable assessment [22, 23]. Preprocessing choices—including channel selection, filtering, and artifact removal—introduce further variability and can bias results when hyperparameters are tuned on overlapping data [19, 24]. Although nested cross‑validation is widely used in other domains for ensembling [19] and hyperparameter optimization [25], it can be applied in PD detection to quantify variability across patient‑level hold‑out sets and channel configurations. In heterogeneous settings with rare events, neglecting such stability checks risks overconfident claims of deployment readiness [21]. Finally, no existing PD studies attempt non‑harmonized cross‑dataset validation (e.g., Iowa vs. San Diego), leaving open whether models trained on one population can generalize to others with different acquisition protocols or demographics.
DL approaches have also been explored for PD detection, but many inherit the same evaluation flaws. For instance, [26] applied convolutional neural networks to EEG recordings of patients and controls with ambiguous enforcement of subject‑independent splits, leaving the reported performance difficult to interpret. Likewise, [13] presented hybrid convolutional models with seemingly strong results, yet overlapping patient data across folds assuredly inflated accuracy estimates. Comparable practices appear in other studies [16, 27, 28, 21], where non‑stratified cross‑validation or single‑block testing at the population level introduce systematic bias. Collectively, these works illustrate how even sophisticated methods can yield misleading outcomes when evaluation fails to guard against leakage and overfitting.
Figure 2: Suggested Workflow for EEG Channel Selection Modeling: Subjects are represented by raw EEG time‑series, which are transformed into spectrograms via short‑time Fourier transform. An inner cross‑validation loop trains a model to estimate per‑channel performance, retaining only the most informative electrodes. These selected channels are reintegrated into the outer training and validation framework, where the final model is trained. Predictions are then aggregated into a single patient‑level probability, providing a concise outcome measure.
A more recent example is the GAN‑Enhanced Generalizable Model for EEG‑Based Detection of Parkinson’s Disease (GEPD) [29], which reported roughly 85% accuracy using generative augmentation. While advancing cross‑dataset generalization, its reliance on a single 90/10 split and train‑on‑Iowa/test‑on‑UNM configuration—without patient‑level aggregation—raises concerns. Similar protocols, including the reverse train/test setup [16, 21], risk tuning models to population‑specific artifacts, producing unstable and inflated performance estimates. Furthermore, GEPD discarding non‑matching EEG channels across datasets limits clinical applicability. More broadly, generative augmentation itself can amplify inherent biases if synthetic samples reproduce population‑specific artifacts rather than disease‑relevant features, underscoring the need for careful validation [30].
In contrast, our design emphasizes rigorous evaluation by integrating three public datasets without discarding viable channels, applying five‑fold stratified cross‑validation with patient grouping, and reporting outcomes at the patient level. These protocols mitigate overfitting to dataset artifacts and yield more reliable estimates of generalization, a critical requirement for biomedical applications. Our approach also advances interpretability, where prior studies treat neural networks as black boxes, we provide band‑level explanations of EEG features, Illuminating which spectral components drive predictions. Thus, methodological robustness and transparent feature attribution demonstrates that competitive performance can be achieved without data leakage or singular validations, while also enhancing trust and clinical relevance.
III Proposed Nested Cross-Validation Architecture
Our framework integrates preprocessing, windowing, and feature extraction with a nested cross-validation design that enforces patient-level separation, supports principled channel selection, and, when desired, incorporates CNN-based band alignment for interpretability. The overall workflow is illustrated in Figure 2 and narrated throughout this section.
III-A Preprocessing
All preprocessing steps are intended to normalize EEG signals within target frequency bands for our nested cross-validation framework. Let $\mathbf{s}_{m}(t)$ denote the raw EEG signal for patient $p_{m}$, sampled at original frequency $f_{\text{orig}}$. Each signal was resampled to a common rate of $f_{s}=64~\text{Hz}$ to preserve oscillatory activity in the Beta (13133030 Hz) and Theta (4488 Hz) bands. The resampled signal $\tilde{\mathbf{s}}_{m}(t)$ was standardized to zero mean and unit variance, $\mathbf{z}_{m}(t)=\frac{\tilde{\mathbf{s}}_{m}(t)-\mu_{m}}{\sigma_{m}},$ where $\mu_{m}$ and $\sigma_{m}$ are the mean and standard deviation of $\tilde{\mathbf{s}}_{m}(t)$. Each standardized trace was then normalized to unit amplitude, $\hat{\mathbf{s}}_{m}(t)=\frac{\mathbf{z}_{m}(t)}{\max_{t}|\mathbf{z}_{m}(t)|}.$ For each subject $p_{m}$, the aligned EEG arrays $\hat{\mathbf{s}}_{m}(t)$ were stored in a DataFrame with metadata (ID, session, diagnosis, file name). Thus, ensuring reproducibility, subject-level traceability, and partitioning within outer cross-validation folds, preventing leakage from the outset.
To ensure comparability across datasets, we incorporated a channel harmonization step. Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{C}\}$ denote the full set of electrodes, where $|C|$ corresponds to the number of channels in the Iowa dataset. Channels were first partitioned into anatomical regions, $\mathcal{C}=\mathcal{C}_{\text{frontal}}\cup\mathcal{C}_{\text{fronto‑central}}\cup\mathcal{C}_{\text{central}}\cup\cdots\cup\mathcal{C}_{\text{occipital}},$ with each subset $\mathcal{C}_{r}$ preserving spatial contiguity. Within each region, electrodes were ordered according to standard EEG conventions:
[…]
After ordering, each channel was assigned a unique integer index $\pi:\mathcal{C}\rightarrow\{0,1,\dots,C-1\},$ such that $\pi(c)$ reflects its position in the harmonized left‑to‑right, anterior‑to‑posterior sequence. This mapping provided a consistent spatial template across datasets, enabling reproducible alignment of electrodes to model inputs and ensuring that subsequent channel selection operated on harmonized, regionally structured arrays. To accommodate datasets with incomplete coverage, any channel $c\in\mathcal{C}$ absent from a given recording was zero-padded at its corresponding index $\pi(c)$. This ensured that all input arrays conformed to the full Iowa-based template, preserving spatial structure while maintaining compatibility across heterogeneous datasets.
III-B Cross-Validation Schema
As illustrated in Figure 2, we adopt a patient‑stratified nested cross‑validation design where the outer loop estimates generalization to unseen patients and the inner loop supports model selection and channel pruning. Both loops use StratifiedGroupKFold with patient identifiers from our DataFrame’s metadata as grouping variables, ensuring all samples $\mathcal{W}_{m}$ from a patient $p_{m}$ are confined to either training or testing folds. This structure prevents patient and temporal leakage while preserving class balance, whereby nesting hyperparameter tuning and channel selection within the inner loop maintains unbiased outer evaluations.
III-C Windowing and Feature Extraction
To define the data segments used in the cross‑validation schema, we employed a windowing strategy tailored to split the ‘Full EEG Signal Stack’ of each patient into channel separated spectrograms, as depicted in Figure 2, which vary in length and recording conditions. With sampling rate $f_{s}=64~\text{Hz}$, the outer window length was set to $L_{w}=16{,}384~\text{samples},$ corresponding to $\frac{16{,}384}{f_{s}}\approx 256~\text{s}.$ Shorter signals were zero‑padded, and longer signals segmented into overlapping windows with hop size $H_{w}=\frac{L_{w}}{4}=4{,}096~\text{samples}.$ Within each window, we computed the short‑time Fourier transform (STFT). For a discrete‑time signal $\mathbf{w}[n]\in\mathbb{R}^{T}$, the STFT is
$\mathbf{STFT}(i,f)=\sum_{k=0}^{L-1}\mathbf{w}[bi+k]\cdot\mathbf{x}[k]\cdot e^{-j2\pi kf/L},$ (2)
where $L$ is the analysis window length, $b$ the hop size, $\mathbf{x}[k]$ the window function, $f$ the frequency bin index, and $i$ the time frame index. This yields a complex‑valued matrix of size $N\times L$, with $N=\left\lfloor\frac{T}{b}\right\rfloor,\quad\Delta f=\frac{f_{s}}{L}.$ We used $n_{\text{fft}}=256$ and $h=64$ samples, producing spectrograms with $F=\frac{n_{\text{fft}}}{2}=128\quad\text{frequency bins},\qquad T=\frac{L_{w}}{h}=\frac{16{,}384}{64}=256\quad\text{time bins}.$ Each EEG channel was thus represented as $(F\times T\times 1)=(128\times 256\times 1),$normalized and converted to decibel scale. This procedure expands each patient’s EEG into multiple channel separated spectrograms, increasing training sample size and enabling finer‑grained decisions while maintaining stable computational demands.
III-D Channel Selection
As highlighted in the ‘Inner Loop Channel Selection’ stage of Figure 2, channel selection is embedded within the nested cross‑validation schema to address the high dimensionality of EEG recordings, where $C\in\{32,64\}$ active electrodes capture distinct neural activity. For each outer fold’s training set, spectrograms from all active channels $\{c_{1},c_{2},\dots,c_{C}\}$ are passed through the CNN, and per‑channel accuracies $a_{c,f}$ are recorded on each inner‑fold validation subset $f\in\{1,\dots,F\}$. These fold‑wise accuracies are then accumulated and averaged to obtain a stability‑adjusted score for each channel where only the active channels for each patient are accumulated into their respective average,
$\bar{a}_{c}=\frac{1}{F}\sum_{f=1}^{F}a_{c,f}.$ (3)
Channels are ranked by $\bar{a}_{c}$, and the top subset $\mathcal{C}^{\ast}\subseteq\{c_{1},\dots,c_{C}\}$ is retained for evaluation in the outer fold; channel selection occurs in the inner loop, ensuring dimensionality reduction is guided by cross‑validated evidence rather than outer‑fold outcomes, while in the outer loop the CNN is retrained on the full channel set and the inner loop identifies electrodes most reliable for inference, maintaining strict separation to preserve patient‑level validity.
III-E CNN Band-Level Alignment
As illustrated in the ‘Single Patient Probability Scores’ stage of Figure 2, we optionally employ the Adaptable Resolution Pooling Network (ARP‑N), a convolutional model trained within both inner and outer loops of the pipeline. ARP‑N transforms variable‑sized spectrograms into consistent square feature maps through convolutional blocks, batch normalization, dropout, and specialized pooling, followed by a dense sigmoid output. For each channel $c$, the STFT spectrogram $\mathbf{X}_{c}\in\mathbb{R}^{F\times T}$ with $F=128$ frequency bins and $T=256$ time bins is adaptively pooled to $\mathbf{Z}_{c}\in\mathbb{R}^{8\times 8}$. Along the frequency axis, $F=128\;\;\longrightarrow\;\;F^{\prime}=8,$ thus spanning
$\Delta f^{\prime}=\frac{32~\text{Hz}}{8}=4~\text{Hz}.$ (4)
This resolution aligns directly with canonical EEG bands— [04][0\!-\!4] Hz (Delta), [48][4\!-\!8] Hz (Theta), [812][8\!-\!12] Hz (Alpha), and [1230][12\!-\!30] Hz (Beta, subdivided)—embedding neuroscientific conventions into the architecture itself. In practice, ARP‑N highlights spectral-temporal regions most indicative of Parkinson’s, providing implicit band‑level interpretability while reducing computational cost and improving generalization across heterogeneous datasets.
III-F Patient-Level Evaluation
As shown in Figure 2 under ‘Patient Decision’, evaluation is performed at the patient-level to reflect clinical use cases. Each patient $p_{m}\in\mathcal{P}$ contributes a set of windowed spectrograms $\mathcal{W}_{m}$, and the model produces window‑level probabilities $\hat{y}_{mi}$. These are aggregated by simple averaging,
$\hat{y}_{m}=\frac{1}{|\mathcal{W}_{m}|}\sum_{i=1}^{|\mathcal{W}_{m}|}\hat{y}_{mi},$ (5)
yielding a single prediction per patient. This procedure ensures that performance metrics correspond to generalization across distinct individuals rather than across overlapping signal segments, thereby aligning the evaluation with the outer cross‑validation design and providing clinically meaningful estimates of model utility.
IV Experiments and Results
We evaluated our framework through a series of controlled experiments spanning multiple datasets, metrics, and model configurations. The following subsections detail the datasets used, evaluation criteria, baseline comparisons, and implementation choices, followed by analyses of model performance (Table I), aggregation strategies (Table II), and interpretability via Grad-CAM visualizations (Figure 3).
IV-A Datasets
We combined three independent EEG dataset’s resting state eyes closed subsets to assess robustness and generalizability: (i) Iowa, 64–channel BioSemi ActiveTwo recordings; (ii) University of New Mexico (UNM), multi-session 64–channel EEG from PD patients and controls; and (iii) San Diego, clinical 32–channel EEG representing hardware variability. All signals were aligned to our canonical montage set forth in III-A, with missing channels zero-filled to preserve indexing. This harmonization enabled cross-dataset evaluation without architectural modification.
IV-B Evaluation Metrics
With the dataset approximately balanced at the patient level, we report accuracy, area under the ROC curve (AUC), precision, and recall without adjustments. These metrics capture overall correctness, threshold discrimination, and the trade‑off between false positives and negatives, providing a robust evaluation under balanced conditions.
IV-C Implementation and Reproducibility
All experiments were implemented in Python with tensorflow/keras. Outputs were stored in DataFrames and logged to CSVs to ensure reproducibility. Training employed Adamax with an exponential learning‑rate schedule, early stopping, and checkpointing. Crucially, reproducibility is anchored by the nested cross‑validation channel selection routine (Algorithm 1), which formalizes how channels are chosen and evaluated across folds. Readers are referred to Algorithm 1 for the full specification.
Algorithm 1 Nested Cross-Validation Channel Selection
1:Full dataset $X_{data}$, labels $y_{data}$, patient groups $G_{data}$, channel indices $C_{data}$, parameters $\theta$, number of inner folds $K$, shared memory slots $L$
2:Best channel subset $\mathcal{C}^{*}$, performance across folds
3:Initialize results list $\mathcal{R}\leftarrow\emptyset$
4:Initialize channel score dictionary $\mathcal{S}[ch]\leftarrow\emptyset$
5:for all outer fold $(X_{train},X_{test})$ do
6:  Store data into shared memory slots $L$
7:  Initialize inner cross-validation with $K$ splits
8:  for $k\leftarrow 1$ to $K$ do
9:   Split into inner-train and inner-validation sets:
10:   $(X^{(k)}_{tr},y^{(k)}_{tr},G^{(k)}_{tr},C^{(k)}_{tr})$
11:   $(X^{(k)}_{val},y^{(k)}_{val},G^{(k)}_{val},C^{(k)}_{val})$
12:   Store data into shared memory slots $L$
13:   $\theta\leftarrow\texttt{runTrain}(L,\text{inner}=True)$
14:   $M\leftarrow\texttt{runChan}(L,\theta)$
15:   for all channel chch in $M$ do
16:     Append accuracy $M[\text{Acc}][ch]$ to $\mathcal{S}[ch]$
17:   end for
18:  end for
19:  Compute channel accuracy: $\bar{a}[ch]\leftarrow\text{mean}(\mathcal{S}[ch])$
20:  $\mathcal{C}^{*}\leftarrow\text{Top-}m\text{ channels by }\bar{a}[ch]$
21:  Store $\mathcal{C}^{*}$ in $L$
22:  $\theta\leftarrow\texttt{runTrain}(L,\text{inner}=False)$
23:  Record results restricted to $(\theta,\mathcal{C}^{*})$ into $\mathcal{R}$
24:end for
25:Aggregate results across folds into dataframe $D$
26:Compute averages per configuration: $\bar{D}\leftarrow\text{mean}(D)$
27:return $\mathcal{C}^{*}$, $\bar{D}$
IV-D Baselines
Table I summarizes performance across three evaluation paradigms. The first, No Stratification, trains without patient‑level separation, permitting temporal and subject leakage; we also include results from Lee et al. [13] and Wu et al. [21] , who used a comparable setup. The second, Single‑Population Blocking, reflects models optimized for a single site or cohort; our I‑4‑C model is shown alongside Anjum et al. [16] and Zhang et al. [29], both evaluated under similar conditions. The third, Stratified Cross‑Validation, is our proposed framework, enforcing patient‑level separation across folds. Here, the All‑Channel model uses all electrodes, while reduced‑channel variants (1‑, 2‑, 4‑, 8‑, 16‑Channel) demonstrate principled channel selection. These comparisons reveal inflated performance under non‑stratified or population‑blocked designs and more clinically relevant estimates with stratified cross‑validation.
IV-E Comparative Analysis
The first paradigm, No Stratification, achieved the highest apparent performance in our experiments (Acc = .985, AUC = .986, Prec = .995, Rec = .978), with Lee et al. [13] and Wu et al. [21] reporting better metrics. However, these results reflect data leakage rather than genuine generalization since patient overlap across folds inflates performance, along with near‑zero variance that further signals overfitting. While the other architecture slightly outperforms ours under flawed evaluation, such comparisons are not meaningful: without patient‑level separation, performance cannot be interpreted as evidence of clinical utility.
Under the Single‑Population Blocking paradigm, our 4‑Channel model decisively outperformed prior approaches. In the Iowa‑blocked evaluation, it achieved Acc = .929, AUC = .929, Prec = .875, and Rec = 1.00—substantially higher than the mid‑.80s accuracies reported by Anjum et al. [16], Zhang et al. [29], and Wu et al. [21] under comparable blocked designs. This margin underscores the strength of our feature extraction and channel selection strategy, even when models are tuned to a single dataset. However, such blocked evaluations can be misleading: by restricting training and testing to the same population, they risk inflating performance through site‑specific artifacts rather than genuine disease‑relevant features. Although population block testing offers a partial safeguard, it remains narrow, as models may still overfit to the unique characteristics of a single cohort. These limitations highlight why cross‑validation with patient‑level separation is essential for establishing true generalizability and clinical validity.
Our framework is the first to provide clinically realistic estimates of generalization on heterogeneous EEG channel datasets by enforcing patient‑level separation through Stratified Cross‑Validation. Within this rigorous setting, the 4‑Channel configuration delivered the strongest overall trade‑off, achieving Acc = .806 $\pm$ .092 and Rec = .801 $\pm$ .134 while maintaining stable precision (.785 $\pm$ .159). The 2‑Channel model offered balanced but slightly weaker performance (Acc = .774, AUC = .797), whereas the 8‑Channel variant favored precision (.819 $\pm$ .108) at the expense of recall (.706 $\pm$ .157). By contrast, the 16‑Channel model underperformed across all metrics, suggesting that excessive electrodes introduce redundancy or noise. These results demonstrate not only the robustness of our channel selection strategy but also the importance of evaluating models under stratified, patient‑level cross‑validation, where inflated performance from site‑specific artifacts is minimized and true generalization can be assessed.
Taken together, these results reveal that inflated scores often stem from temporal or patient leakage rather than true predictive ability, and that best‑case performance within a single dataset can mask poor generalization under heterogeneous conditions. By contrast, our stratified framework enforces patient‑level separation, yielding more conservative but clinically meaningful estimates. Moderate channel configurations consistently provided the most reliable balance, whereas both minimal and excessive channel counts degraded performance, underscoring the value of principled selection in the inner loop. Overall, findings highlight the need of stratification and the unique contribution of our framework in aligning evaluation with clinical deployment.
IV-F Patient-Level Aggregation Ablation
Table II presents an ablation of aggregation strategies for patient-level decisions in our nested cross-validation framework. Across all configurations, Mean and Median aggregation consistently produced the most balanced outcomes, reinforcing their reliability for patient‑level evaluation. The 4‑Channel models performed best overall, with both rules yielding strong accuracy and recall, while 2‑Channel models also improved on the baseline. Other strategies showed less stability: Majority performed competitively but leaned toward higher precision, Geometric Mean (GMean) was slightly weaker yet serviceable, and Max and Min proved unreliable—either overemphasizing precision at the cost of recall or collapsing under outliers. Importantly, the Iowa Blocked Test followed this same general trend: Mean, Median, and Majority aggregation produced the strongest and most stable results, while Min and Max underperformed.
Overall, these findings show that aggregation is not a secondary detail but a central design choice that directly shapes clinical validity. Robust rules such as Mean and Median not only stabilize fold‑to‑fold variance but also preserve the sensitivity–specificity balance needed for screening contexts. Conversely, unstable rules amplify the same risks seen with poor stratification or excessive channel counts; namely, poor generalization. Thus, careful aggregation complements stratified cross‑validation and principled channel selection, forming our framework for reliable patient‑level evaluation.
TABLE I: Performance comparison under different stratification strategies and channel subsets.
Model Acc $\sigma$ AUC $\sigma$ Prec $\sigma$ Rec $\sigma$
No Stratification
Wu [21] 1.00 NA 1.00 NA 1.00 NA 1.00 NA
Lee [13] .992 NA .992 NA .989 NA .994 NA
Ours .985 0 .986 .001 .995 .008 .978 .008
Single-Population Blocking
I-4C (Ours) .929 NA 0.929 NA .875 NA 1.00 NA
Anjum [16] .857 NA 0.852 NA .857 NA .857 NA
Wu [21] .857 NA .841 NA .916 NA .786 NA
Zhang [29] .843 NA NA NA .840 NA .840 NA
Stratified Cross-Validation
All‑Channel .715 .070 .738 .102 .711 .091 .759 .131
1-Channel .752 .124 .797 .130 .777 .158 .714 .136
2-Channel .774 .107 .797 .104 .785 .159 .742 .125
4-Channel .806 .092 .799 .067 .785 .159 .801 .134
8-Channel .767 .117 .752 .154 .819 .108 .706 .157
16-Channel .748 .092 .753 .143 .793 .112 .703 .154
TABLE II: Ablation of aggregation strategies for patient-level decision making under different channel configurations.
Method Acc $\sigma$ AUC $\sigma$ Prec $\sigma$ Rec $\sigma$
Single-Population Blocking
I-4C GMean .821 NA .913 NA .846 NA .786 NA
I-4C Majority .929 NA .944 NA .875 NA 1.00 NA
I-4C Max .607 NA .957 NA .560 NA 1.00 NA
I-4C Mean .893 NA .944 NA .824 NA 1.00 NA
I-4C Median .929 NA .944 NA .875 NA 1.00 NA
I-4C Min .571 NA .832 NA 1.00 NA .143 NA
Stratified Cross-Validation
All GMean .685 .103 .734 .095 .608 .153 .777 .143
All Majority .715 .070 .738 .102 .711 .091 .759 .131
All Max .633 .135 .731 .098 .867 .130 .615 .145
All Mean .707 .084 .738 .102 .693 .115 .751 .147
All Median .715 .070 .734 .101 .711 .091 .759 .131
All Min .573 .147 .694 .116 .310 .230 .790 .262
2C GMean .785 .092 .806 .086 .785 .159 .766 .130
2C Majority .766 .106 .797 .104 .807 .123 .721 .130
2C Max .766 .106 .814 .101 .807 .123 .721 .130
2C Mean .774 .107 .797 .104 .785 .159 .742 .125
2C Median .774 .107 .801 .104 .785 .159 .742 .125
2C Min .749 .091 .806 .087 .721 .187 .750 .126
4C GMean .756 .091 .796 .052 .689 .165 .777 .160
4C Majority .801 .084 .799 .067 .829 .099 .762 .141
4C Max .765 .100 .816 .072 .829 .099 .711 .129
4C Mean .806 .092 .799 .067 .785 .159 .801 .134
4C Median .797 .082 .826 .052 .807 .123 .770 .133
4C Min .717 .095 .748 .074 .608 .183 .750 .199
Figure 3: Grad‑CAM overlays for Patient 123 highlight consistent band‑level focus. Each row shows one EEG channel (32, 41, 54, 55), progressing from spectrogram (yellow–green) to Grad‑CAM heatmap (red) to overlay. The model emphasizes recurrent theta activity (4–8 Hz), with secondary attention in alpha (8–12 Hz) and occasional beta (12–30 Hz). These theta–alpha patterns broaden and narrow rhythmically, producing a pulsing appearance distinct from noise or artifacts. While edge effects (e.g., aliasing, zero‑padding) sometimes attract attention, the dominant focus remains on physiologically plausible oscillations, underscoring both the strengths and limits of convolutional interpretability
IV-G Grad-CAM Interpretability Visualization
Grad-CAM overlays in Figure 3 revealed that the network consistently emphasized the theta band (4–8 Hz) across channels, with secondary contributions from alpha (8–12 Hz) and beta (12–30 Hz) ranges. In several cases, the model highlighted recurrent theta–alpha structures with rhythmic fluctuations in spectral bandwidth, producing a pulsing appearance distinct from artifacts. While the CNN occasionally attended to processing artifacts (e.g., zero-padding or aliasing), the dominant focus remained within physiologically meaningful bands; therefore aligning with canonical oscillatory rhythms suggesting learned representations are discriminative and neurophysiologically interpretable. The frequency-specific spatially distributed activations across channels further indicate that the network captured cross-channel dependencies rather than isolated channel effects.
This pattern aligns with prior EEG studies in PD, which report consistent alterations in theta and beta. For example, increased frontal theta activity and impaired beta desynchronization during lower-limb movement have been linked to motor deficits and freezing of gait (FOG) in PD patients [31]. Complementary work using linear predictive coding shows PD patients exhibit not only theta and beta abnormalities but also reduced alpha power, distinguishing them from healthy controls [16]. Thus, our Grad-CAM results suggest that ARP‑N is aligned with clinical research. By discretizing the 0–32 Hz spectrum into 4 Hz bins, the network is structurally matched to canonical EEG bands, encouraging band‑level representations that Grad-CAM makes visible, thereby strengthening confidence in interpretability.
V Conclusion
We introduced a nested cross-validation framework for early detection of PD from EEG, designed as a reproducible and generalizable response to common methodological pitfalls. By harmonizing three independent datasets (Iowa, UNM, San Diego) into a consistent montage, we established a foundation for cross‑site evaluation. Patient stratification in the outer loop prevented data leakage, while multi‑layered windowing preserved temporal structure and enabled uniform processing of heterogeneous recordings. The inner loop enabled principled channel selection, reducing dimensionality and improving generalization by emphasizing the most informative electrodes. Patient‑level aggregation, particularly mean and median, further reinforced clinical reliability. Collectively, these components show nested cross-validation as a principled model‑agnostic solution to leakage, dimensionality, and heterogeneity challenges that have historically inflated performance. Beyond PD, our architecture offers a reproducible framework for EEG and ML studies emphasizing scientific rigor, transparency, and clinical applicability.
Acknowledgments
The author acknowledges Microsoft Copilot for assistance with grammar and structural edits, contextual extrapolation in the results section, and figure preparation with extensive human input [32]. All core ideas, analyses, and interpretations are the author’s own based on previous literature and experimental analysis, as cited in the manuscript.

### v2

@@
 Abstract
-The early detection of Parkinson’s disease remains a critical challenge in clinical neuroscience, with electroencephalography offering a non‑invasive and scalable pathway toward population‑level screening.
+The early detection of Parkinson’s disease remains a critical challenge in clinical neuroscience, with electroencephalography (EEG) offering a non‑invasive and scalable pathway toward population‑level screening.
 While machine learning has shown promise in this domain, many reported results suffer from methodological flaws, most notably patient‑level data leakage, inflating performance estimates and limiting clinical translation.
@@
 I Introduction
-Parkinson’s disease (PD) is the second most common neurodegenerative disorder, affecting over 10 million people worldwide and imposing a growing socioeconomic burden [1].
-Its progressive course—marked by motor dysfunction, cognitive decline, and reduced quality of life—makes early detection and intervention a critical priority  [2, 3].
-Yet current clinical diagnosis is largely based on overt motor symptoms, which typically emerge only after substantial dopaminergic neuron loss [2, 3].
-This delay highlights the importance of identifying prodromal biomarkers, such as REM sleep behavior disorder, olfactory impairment, and subtle motor irregularities, which often precede clinical diagnosis by years [2].
-Detecting early PD could enable interventions when therapies are effective [2].
-Advances in biomedical signal acquisition have created new opportunities for non‑invasive monitoring of these prodromal and progressive symptoms [4].
-Among available modalities, electroencephalography (EEG) is particularly promising: it is portable, relatively inexpensive, and offers millisecond‑level temporal resolution of neural dynamics [5].
-Compared to imaging techniques such as MRI or PET, which provide high spatial resolution but are costly and less accessible, EEG enables scalable, repeated assessments that are well‑suited for longitudinal monitoring [5, 6].
-This combination of accessibility and temporal precision positions EEG as a strong candidate for early PD biomarker discovery.
-Figure 1: The nested cross‑validation architecture addresses the challenge of high‑dimensional EEG data by starting with all channels and systematically reducing to the most informative.
-The inner loop selects channels contributing most to inference, while the outer loop validates this choice without bias.
-This principled “many‑to‑best” reduction is dataset‑, channel‑, and model‑agnostic, yielding results that are more efficient, generalizable, and interpretable.
+Parkinson’s disease (PD) is the second most common neurodegenerative disorder, affecting over 10 million people worldwide and imposing a substantial socioeconomic burden, yet clinical diagnosis typically relies on overt motor symptoms that emerge only after significant neuronal loss [1, 2, 3].
+This delay underscores the need to identify prodromal biomarkers—such as REM sleep behavior disorder and subtle motor changes—that may precede diagnosis by years and enable earlier intervention [2].
+Recent advances in biomedical signal acquisition have made such early monitoring increasingly feasible, with EEG emerging as a particularly promising modality for non‑invasive assessment of prodromal and progressive PD symptoms due to its portability, low cost, and high temporal resolution [4, 5].
+Unlike MRI or PET, EEG enables scalable, repeated assessments, supporting longitudinal monitoring and biomarker discovery [5, 6].
+Figure 1: The nested cross‑validation architecture is applied to high‑dimensional EEG by training models on all channels, while inner‑loop validation ranks channel relevance and defines the optimal test‑time architecture.
+The outer loop evaluates this selected configuration, enabling a dataset‑, channel‑, and model‑agnostic “many‑to‑best” reduction that supports efficient, generalizable, and interpretable inference.
 In parallel, machine learning (ML) and deep learning (DL) methods show strong potential for extracting subtle biomarkers from high‑dimensional biomedical data [7].
-A wide range of neural architectures have been explored for biosignal analysis: recurrent neural networks (RNNs) and long short‑term memory (LSTM) models capture temporal dependencies in sequential data [8]; transformer‑based models leverage self‑attention to model long‑range interactions across time and frequency; and graph neural networks (GNNs) exploit the spatial topology of EEG electrode layouts to learn structured representations [9].
+A wide range of neural architectures have been explored for biosignal analysis, including temporal recurrent models, self‑attention based transformer models, and graph neural networks that exploit EEG electrode layouts [8, 9].
 Despite these advances, convolutional neural networks (CNNs) remain a principled starting point for EEG analysis.
-When applied to short‑time Fourier transform (STFT) spectrograms, CNNs exploit partial shift‑invariance that complements the local stationarity of the STFT, enabling robust feature extraction without handcrafted preprocessing [10, 11, 12].
+When applied to short‑time Fourier transform (STFT) spectrograms, CNNs exploit partial shift‑invariance aligned with the local stationarity of the STFT [10, 11, 12].
 This synergy has consistently delivered state‑of‑the‑art performance across diverse signal domains, including PD detection, where CNN‑based or hybrid pipelines frequently report accuracies exceeding 95% [13].
-Moreover, CNN architectures are naturally extensible to multimodal fusion: EEG inputs can be paired with bio-signal sensor data through parallel branches converging in shared layers, allowing integrated biomarker discovery without redesigning the core model [14].
-Many reported successes in EEG‑based PD detection are undermined by methodological flaws, most notably data leakage and dataset heterogeneity.
-Leakage arises in two primary forms: temporal leakage, where overlapping windows from the same recording are split across training and test sets, and subject leakage, where data from the same patient appears in both folds.
-In both cases, models risk memorizing idiosyncratic traits rather than learning disease‑relevant features, inflating performance and eroding clinical validity [15].
-Surveys of PD classification studies further highlight that EEG research often employs ambiguous patient‑level stratification, if it is applied at all [16].
-Population block testing offers a partial safeguard, yet its scope is limited, as models may still overfit to the unique characteristics of a single cohort.
-These concerns are compounded by the heterogeneity of public EEG datasets: channel counts range from high‑density 64‑channel caps to sparse consumer devices, and acquisition hardware differs in sampling rate, electrode placement, and preprocessing pipelines.
-Methods tuned to one configuration often fail to transfer across others, creating a significant reproducibility barrier.
-Collectively, these challenges underscore the need for architectures that prevent leakage, generalize across populations, and robustly reduce high‑dimensional signals to the most informative subset, independent of dataset or recording setup.
-To address these challenges, we propose a nested cross-validation framework for robust EEG analysis (Figure 1).
-The outer loop enforces patient-level stratification to eliminate leakage, while the inner loop selects the most informative channels, reducing dimensionality without sacrificing clinically relevant signals.
-A multi-layered windowing strategy integrates recordings of varying length, sampling rate, and hardware, enabling consistent evaluation across heterogeneous datasets.
-By design, the framework is dataset, channel, and model-agnostic, positioning it as a blueprint for reproducible EEG-based biomarker discovery.
-Together, these components yield generalizable performance estimates and clinically meaningful early PD detection results.
-Our contributions are threefold:
-  1. 1.
-    A model‑agnostic nested cross‑validation framework with stratification, windowing, and inner‑loop channel selection to prevent leakage, harmonize datasets, and reduce dimensionality.
-  2. 2.
-    Empirical evidence that patient‑level leakage and narrow cohort blocking inflate EEG‑based PD detection performance and undermine clinical validity.
-  3. 3.
-    Demonstration that learned representations align with canonical EEG frequency bands, yielding interpretable models that achieve 0.806 accuracy across datasets.
-Looking forward, this framework provides a foundation for accelerating clinical translation: it enables reproducible biomarker validation across heterogeneous cohorts, supports integration with multimodal data streams such as EMG and wearable sensors, and lays the groundwork for scalable digital health pipelines capable of monitoring neurodegenerative disease progression in real-world settings.
+Moreover, CNN architectures are extensible to multimodal fusion, allowing EEG to be integrated with complementary biosignals through parallel branches [14].
+Many reported successes in EEG‑based PD detection are undermined by methodological issues, particularly data leakage and dataset heterogeneity.
+Leakage from overlapping temporal windows or subject overlap allows models to exploit patient‑specific idiosyncrasies rather than disease‑relevant features [15], and population‑level splits provide only partial protection against cohort‑specific overfitting [16].
+Combined with heterogeneity in EEG channel counts, hardware, and preprocessing, these factors limit reproducibility and cross‑dataset transferability, motivating evaluation frameworks that explicitly prevent leakage and assess generalization across diverse populations.
+To address these challenges, we propose a nested cross‑validation framework for robust EEG analysis (Figure 1) that enforces patient‑level stratification in the outer loop to eliminate leakage, performs channel selection in the inner loop to reduce dimensionality without discarding clinically relevant signals, and integrates a multi‑layered windowing strategy to accommodate recordings with varying length, sampling rate, and hardware.
+By design, the framework is dataset‑, channel‑, and model‑agnostic, positioning it as a reproducible blueprint for EEG‑based biomarker discovery.
+Accordingly, our contributions are threefold: (i) a model‑agnostic, stratified nested cross‑validation framework with integrated channel selection to prevent leakage, harmonize heterogeneous datasets, and reduce dimensionality; (ii) empirical evidence that patient‑level leakage and narrow cohort blocking inflate reported EEG‑based PD detection performance and undermine clinical validity; and (iii) demonstration that learned representations align with canonical EEG frequency bands, yielding interpretable models that achieve 80.6% accuracy across datasets.
+Collectively, this framework supports clinical translation by enabling reproducible biomarker validation across diverse cohorts, facilitating integration with multimodal data (e.g., EMG and wearable sensors), and supporting scalable digital health pipelines for monitoring neurodegenerative disease progression in real‑world settings using ML.
 II Related Work and Methodological Pitfalls
-A growing body of research has applied ML techniques to the early detection of PD using EEG, EMG, and other biosignal modalities.
-Several studies have reported high classification accuracies, often exceeding 85%, suggesting strong potential for clinical translation [17, 18].
-However, closer inspection reveals that many of these works rely on evaluation protocols that do not stratify data at the patient level.
-Instead, temporal windows from the same subject are distributed across both training and test sets, creating a form of data leakage.
-This allows models to exploit subject‑specific idiosyncrasies rather than learning generalizable disease markers, thereby inflating performance estimates [15, 19, 20, 21].
-Beyond this central issue, several additional pitfalls recur across the literature.
-The small sample sizes typical of PD datasets magnify overfitting risks, as even a few misallocated subjects can distort reported outcomes [20].
-Reported performance is also often reduced to accuracy, which obscures clinically relevant trade‑offs; complementary metrics such as sensitivity, specificity, ROC‑AUC, and calibration curves provide a more reliable assessment [22, 23].
-Preprocessing choices—including channel selection, filtering, and artifact removal—introduce further variability and can bias results when hyperparameters are tuned on overlapping data [19, 24].
-Although nested cross‑validation is widely used in other domains for ensembling [19] and hyperparameter optimization [25], it can be applied in PD detection to quantify variability across patient‑level hold‑out sets and channel configurations.
-In heterogeneous settings with rare events, neglecting such stability checks risks overconfident claims of deployment readiness [21].
-Finally, no existing PD studies attempt non‑harmonized cross‑dataset validation (e.g., Iowa vs.
-San Diego), leaving open whether models trained on one population can generalize to others with different acquisition protocols or demographics.
-DL approaches have also been explored for PD detection, but many inherit the same evaluation flaws.
-For instance, [26] applied convolutional neural networks to EEG recordings of patients and controls with ambiguous enforcement of subject‑independent splits, leaving the reported performance difficult to interpret.
-Likewise, [13] presented hybrid convolutional models with seemingly strong results, yet overlapping patient data across folds assuredly inflated accuracy estimates.
-Comparable practices appear in other studies [16, 27, 28, 21], where non‑stratified cross‑validation or single‑block testing at the population level introduce systematic bias.
-Collectively, these works illustrate how even sophisticated methods can yield misleading outcomes when evaluation fails to guard against leakage and overfitting.
-Figure 2: Suggested Workflow for EEG Channel Selection Modeling: Subjects are represented by raw EEG time‑series, which are transformed into spectrograms via short‑time Fourier transform.
-An inner cross‑validation loop trains a model to estimate per‑channel performance, retaining only the most informative electrodes.
-These selected channels are reintegrated into the outer training and validation framework, where the final model is trained.
-Predictions are then aggregated into a single patient‑level probability, providing a concise outcome measure.
-A more recent example is the GAN‑Enhanced Generalizable Model for EEG‑Based Detection of Parkinson’s Disease (GEPD) [29], which reported roughly 85% accuracy using generative augmentation.
-While advancing cross‑dataset generalization, its reliance on a single 90/10 split and train‑on‑Iowa/test‑on‑UNM configuration—without patient‑level aggregation—raises concerns.
-Similar protocols, including the reverse train/test setup [16, 21], risk tuning models to population‑specific artifacts, producing unstable and inflated performance estimates.
-Furthermore, GEPD discarding non‑matching EEG channels across datasets limits clinical applicability.
-More broadly, generative augmentation itself can amplify inherent biases if synthetic samples reproduce population‑specific artifacts rather than disease‑relevant features, underscoring the need for careful validation [30].
-In contrast, our design emphasizes rigorous evaluation by integrating three public datasets without discarding viable channels, applying five‑fold stratified cross‑validation with patient grouping, and reporting outcomes at the patient level.
-These protocols mitigate overfitting to dataset artifacts and yield more reliable estimates of generalization, a critical requirement for biomedical applications.
-Our approach also advances interpretability, where prior studies treat neural networks as black boxes, we provide band‑level explanations of EEG features, Illuminating which spectral components drive predictions.
-Thus, methodological robustness and transparent feature attribution demonstrates that competitive performance can be achieved without data leakage or singular validations, while also enhancing trust and clinical relevance.
+A growing body of work applies ML techniques to early PD detection using EEG, EMG, and other biosignal modalities, with several studies reporting classification accuracies exceeding 85%, suggesting strong translational potential [17, 18].
+However, many of these results rely on evaluation protocols that fail to enforce patient‑level stratification, allowing temporal windows from the same subject to appear in both training and test sets and introducing data leakage.
+This enables models to exploit subject‑specific idiosyncrasies rather than learning generalizable disease markers, inflating reported performance estimates [15, 19, 20, 21].
+Beyond data leakage, EEG‑based PD detection is limited by small sample sizes, accuracy‑only reporting, and preprocessing‑induced bias.
+Minor subject misallocation can substantially inflate performance [20], while reliance on accuracy obscures clinically relevant trade‑offs [22, 23].
+Preprocessing choices further bias results when hyperparameters are tuned on overlapping data [19, 24].
+Although nested cross‑validation is standard for ensembling and optimization [19, 25], its use in PD detection would enable stability assessment across patient‑level splits and channel configurations, reducing overconfidence [21].
+Thusly, no existing studies evaluate non‑aligned cross‑dataset generalization, leaving cross‑protocol transferability unexplored.
+Figure 2: Workflow for EEG channel selection: raw EEG is transformed into spectrograms, channels are ranked via inner‑loop cross‑validation, selected electrodes are evaluated in the outer loop, and window‑level predictions are aggregated into a single patient‑level probability.
+Deep learning approaches for EEG‑based PD detection often inherit evaluation weaknesses, including non‑independent subject splits that inflate reported performance [26, 13].
+Similar bias arises from non‑stratified cross‑validation and single‑population testing in later studies [16, 27, 28].
+Although recent work explores cross‑dataset generalization via generative augmentation with reported accuracies near 85% [29], such methods risk amplifying population‑specific artifacts when synthetic data reproduce dataset biases rather than disease‑relevant structure [30].
+Among recent studies, Wu et al.
+[21] represent a methodological advance by explicitly addressing channel selection in EEG‑based PD detection.
+Their approach evaluates individual channels using a single train–validation split on the UNM dataset [16], retaining electrodes that exceed a fixed performance threshold before retraining on the selected subset.
+While this provides insight into channel relevance and improves performance under population‑blocked testing, selection is performed once and remains fixed across runs, without accounting for variability from data partitioning or patient‑level heterogeneity.
+Moreover, the procedure is tightly coupled to a specific dataset and feature representation, limiting robustness across heterogeneous cohorts.
+Our framework embeds channel selection within nested, patient‑stratified cross‑validation, replacing single‑pass, dataset‑specific procedures with stability‑adjusted estimates derived from repeated inner‑fold evaluations.
+All channels are retained during outer‑loop training to preserve shared spatial representations, while inference is restricted to the most reliable electrodes, yielding robust performance without leakage, singular validations, or dataset‑specific tuning and enabling more stringent, clinically relevant assessments of generalization under heterogeneous conditions.
 III Proposed Nested Cross-Validation Architecture
@@
 III-A Preprocessing
-All preprocessing steps are intended to normalize EEG signals within target frequency bands for our nested cross-validation framework.
-Let $\mathbf{s}_{m}(t)$ denote the raw EEG signal for patient $p_{m}$, sampled at original frequency $f_{\text{orig}}$.
-Each signal was resampled to a common rate of $f_{s}=64~\text{Hz}$ to preserve oscillatory activity in the Beta (13133030 Hz) and Theta (4488 Hz) bands.
-The resampled signal $\tilde{\mathbf{s}}_{m}(t)$ was standardized to zero mean and unit variance, $\mathbf{z}_{m}(t)=\frac{\tilde{\mathbf{s}}_{m}(t)-\mu_{m}}{\sigma_{m}},$ where $\mu_{m}$ and $\sigma_{m}$ are the mean and standard deviation of $\tilde{\mathbf{s}}_{m}(t)$.
-Each standardized trace was then normalized to unit amplitude, $\hat{\mathbf{s}}_{m}(t)=\frac{\mathbf{z}_{m}(t)}{\max_{t}|\mathbf{z}_{m}(t)|}.$ For each subject $p_{m}$, the aligned EEG arrays $\hat{\mathbf{s}}_{m}(t)$ were stored in a DataFrame with metadata (ID, session, diagnosis, file name).
-Thus, ensuring reproducibility, subject-level traceability, and partitioning within outer cross-validation folds, preventing leakage from the outset.
-To ensure comparability across datasets, we incorporated a channel harmonization step.
-Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{C}\}$ denote the full set of electrodes, where $|C|$ corresponds to the number of channels in the Iowa dataset.
-Channels were first partitioned into anatomical regions, $\mathcal{C}=\mathcal{C}_{\text{frontal}}\cup\mathcal{C}_{\text{fronto‑central}}\cup\mathcal{C}_{\text{central}}\cup\cdots\cup\mathcal{C}_{\text{occipital}},$ with each subset $\mathcal{C}_{r}$ preserving spatial contiguity.
-Within each region, electrodes were ordered according to standard EEG conventions:
+EEG Signal Preprocessing Steps
+Let $\mathbf{s}_{m}(t)$ denote the raw EEG signal for patient $p_{m}$, sampled at frequency $f_{\text{orig}}$.
+Signals were resampled to $f_{s}=64~\text{Hz}$ to preserve Beta (13133030 Hz) and Theta (4488 Hz) activity, yielding $\tilde{\mathbf{s}}_{m}(t)$.
+Each trace was standardized, $\mathbf{z}_{m}(t)=\frac{\tilde{\mathbf{s}}_{m}(t)-\mu_{m}}{\sigma_{m}},$ and normalized to unit amplitude, $\hat{\mathbf{s}}_{m}(t)=\frac{\mathbf{z}_{m}(t)}{\max_{t}|\mathbf{z}_{m}(t)|}.$ Aligned signals $\hat{\mathbf{s}}_{m}(t)$ were stored with subject metadata to ensure traceability and leakage‑free partitioning across outer folds.
+To ensure cross‑dataset comparability, channels were harmonized to a common spatial template.
+Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{C}\}$ denote the full electrode set, with $|C|$ defined by the Iowa dataset.
+Channels were grouped into contiguous anatomical regions, $\mathcal{C}=\mathcal{C}_{\text{frontal}}\cup\mathcal{C}_{\text{fronto‑central}}\cup\mathcal{C}_{\text{central}}\cup\cdots\cup\mathcal{C}_{\text{occipital}},$ and ordered within each region according to standard EEG conventions:
 […]
-After ordering, each channel was assigned a unique integer index $\pi:\mathcal{C}\rightarrow\{0,1,\dots,C-1\},$ such that $\pi(c)$ reflects its position in the harmonized left‑to‑right, anterior‑to‑posterior sequence.
-This mapping provided a consistent spatial template across datasets, enabling reproducible alignment of electrodes to model inputs and ensuring that subsequent channel selection operated on harmonized, regionally structured arrays.
-To accommodate datasets with incomplete coverage, any channel $c\in\mathcal{C}$ absent from a given recording was zero-padded at its corresponding index $\pi(c)$.
-This ensured that all input arrays conformed to the full Iowa-based template, preserving spatial structure while maintaining compatibility across heterogeneous datasets.
-III-B Cross-Validation Schema
-As illustrated in Figure 2, we adopt a patient‑stratified nested cross‑validation design where the outer loop estimates generalization to unseen patients and the inner loop supports model selection and channel pruning.
-Both loops use StratifiedGroupKFold with patient identifiers from our DataFrame’s metadata as grouping variables, ensuring all samples $\mathcal{W}_{m}$ from a patient $p_{m}$ are confined to either training or testing folds.
-This structure prevents patient and temporal leakage while preserving class balance, whereby nesting hyperparameter tuning and channel selection within the inner loop maintains unbiased outer evaluations.
-III-C Windowing and Feature Extraction
-To define the data segments used in the cross‑validation schema, we employed a windowing strategy tailored to split the ‘Full EEG Signal Stack’ of each patient into channel separated spectrograms, as depicted in Figure 2, which vary in length and recording conditions.
-With sampling rate $f_{s}=64~\text{Hz}$, the outer window length was set to $L_{w}=16{,}384~\text{samples},$ corresponding to $\frac{16{,}384}{f_{s}}\approx 256~\text{s}.$ Shorter signals were zero‑padded, and longer signals segmented into overlapping windows with hop size $H_{w}=\frac{L_{w}}{4}=4{,}096~\text{samples}.$ Within each window, we computed the short‑time Fourier transform (STFT).
-For a discrete‑time signal $\mathbf{w}[n]\in\mathbb{R}^{T}$, the STFT is
+Each channel was assigned an index $\pi:\mathcal{C}\rightarrow\{0,1,\dots,C-1\},$ reflecting its left‑to‑right, anterior‑to‑posterior position.
+Channels absent from a recording were zero‑padded at index $\pi(c)$, ensuring all inputs conform to a consistent spatial structure across heterogeneous datasets.
+Cross‑Validation Schema
+We employ a patient‑stratified nested cross‑validation design in which the outer loop estimates generalization to unseen patients and the inner loop supports channel selection (Figure 2).
+Both loops use StratifiedGroupKFold with patient identifiers as grouping variables, ensuring all samples $\mathcal{W}_{m}$ from a patient $p_{m}$ remain within a single fold.
+This structure prevents patient and temporal leakage, preserves class balance, and maintains unbiased outer‑loop evaluation by nesting channel selection within the inner loop.
+Windowing and Feature Extraction
+To accommodate variable recording lengths and conditions, each patient’s EEG signal was segmented into channel‑wise spectrograms (Figure 2).
+With sampling rate $f_{s}=64~\text{Hz}$, signals were windowed using $L_{w}=16{,}384~\text{samples}\;(\approx 256~\text{s}),$ with shorter recordings zero‑padded and longer recordings segmented using hop size $H_{w}=\frac{L_{w}}{2}=8{,}192~\text{samples}.$ Within each window, the short‑time Fourier transform (STFT) was computed.
+For a discrete‑time signal $\mathbf{w}[n]\in\mathbb{R}^{L_{w}}$,
 $\mathbf{STFT}(i,f)=\sum_{k=0}^{L-1}\mathbf{w}[bi+k]\cdot\mathbf{x}[k]\cdot e^{-j2\pi kf/L},$ (2)
 where $L$ is the analysis window length, $b$ the hop size, $\mathbf{x}[k]$ the window function, $f$ the frequency bin index, and $i$ the time frame index.
-This yields a complex‑valued matrix of size $N\times L$, with $N=\left\lfloor\frac{T}{b}\right\rfloor,\quad\Delta f=\frac{f_{s}}{L}.$ We used $n_{\text{fft}}=256$ and $h=64$ samples, producing spectrograms with $F=\frac{n_{\text{fft}}}{2}=128\quad\text{frequency bins},\qquad T=\frac{L_{w}}{h}=\frac{16{,}384}{64}=256\quad\text{time bins}.$ Each EEG channel was thus represented as $(F\times T\times 1)=(128\times 256\times 1),$normalized and converted to decibel scale.
-This procedure expands each patient’s EEG into multiple channel separated spectrograms, increasing training sample size and enabling finer‑grained decisions while maintaining stable computational demands.
-III-D Channel Selection
-As highlighted in the ‘Inner Loop Channel Selection’ stage of Figure 2, channel selection is embedded within the nested cross‑validation schema to address the high dimensionality of EEG recordings, where $C\in\{32,64\}$ active electrodes capture distinct neural activity.
-For each outer fold’s training set, spectrograms from all active channels $\{c_{1},c_{2},\dots,c_{C}\}$ are passed through the CNN, and per‑channel accuracies $a_{c,f}$ are recorded on each inner‑fold validation subset $f\in\{1,\dots,F\}$.
-These fold‑wise accuracies are then accumulated and averaged to obtain a stability‑adjusted score for each channel where only the active channels for each patient are accumulated into their respective average,
+This yields a complex time–frequency representation indexed by $(i,f)$, with the number of time frames and frequency resolution given by $T=\left\lfloor\frac{L_{w}}{b}\right\rfloor,\quad\Delta f=\frac{f_{s}}{L}.$ Using a discrete FFT implementation with $n_{\text{fft}}=256$ and hop size $h=64$, each channel produces magnitude spectrograms with $F=\frac{n_{\text{fft}}}{2}=128$ frequency bins and $T=\frac{L_{w}}{h}=256$ time frames, stored as $(F\times T\times 1)=(128\times 256\times 1)$ tensors, converted to decibel scale and normalized.
+III-B Inner Loop Channel Selection
+As shown in Figure 2, channel selection is embedded within the nested cross‑validation framework to address the high dimensionality of EEG recordings with $C\in\{32,64\}$ electrodes.
+For each outer‑fold training set, per‑channel accuracies $a_{c,f}$ are computed across inner‑folds $f\in\{1,\dots,F\}$ and averaged to obtain a stability‑adjusted score,
 $\bar{a}_{c}=\frac{1}{F}\sum_{f=1}^{F}a_{c,f}.$ (3)
-Channels are ranked by $\bar{a}_{c}$, and the top subset $\mathcal{C}^{\ast}\subseteq\{c_{1},\dots,c_{C}\}$ is retained for evaluation in the outer fold; channel selection occurs in the inner loop, ensuring dimensionality reduction is guided by cross‑validated evidence rather than outer‑fold outcomes, while in the outer loop the CNN is retrained on the full channel set and the inner loop identifies electrodes most reliable for inference, maintaining strict separation to preserve patient‑level validity.
-III-E CNN Band-Level Alignment
-As illustrated in the ‘Single Patient Probability Scores’ stage of Figure 2, we optionally employ the Adaptable Resolution Pooling Network (ARP‑N), a convolutional model trained within both inner and outer loops of the pipeline.
+Channels are ranked by $\bar{a}_{c}$, and the top subset $\mathcal{C}^{\ast}$ is selected in the inner loop, ensuring dimensionality reduction is guided by cross‑validated evidence while preserving strict separation from outer‑fold evaluation preventing leakage.
+III-C Patient Decision
+Adaptive Resolution Pooling Network (ARP‑N)
+As illustrated in the ‘Single Patient Probability Scores’ stage of Figure 2, we optionally employ the Adaptive Resolution Pooling Network (ARP‑N) [31].
 ARP‑N transforms variable‑sized spectrograms into consistent square feature maps through convolutional blocks, batch normalization, dropout, and specialized pooling, followed by a dense sigmoid output.
@@
 This resolution aligns directly with canonical EEG bands— [04][0\!-\!4] Hz (Delta), [48][4\!-\!8] Hz (Theta), [812][8\!-\!12] Hz (Alpha), and [1230][12\!-\!30] Hz (Beta, subdivided)—embedding neuroscientific conventions into the architecture itself.
-In practice, ARP‑N highlights spectral-temporal regions most indicative of Parkinson’s, providing implicit band‑level interpretability while reducing computational cost and improving generalization across heterogeneous datasets.
-III-F Patient-Level Evaluation
-As shown in Figure 2 under ‘Patient Decision’, evaluation is performed at the patient-level to reflect clinical use cases.
-Each patient $p_{m}\in\mathcal{P}$ contributes a set of windowed spectrograms $\mathcal{W}_{m}$, and the model produces window‑level probabilities $\hat{y}_{mi}$.
-These are aggregated by simple averaging,
+In practice, ARP‑N highlights spectral-temporal regions most indicative of Parkinson’s Disease, providing implicit band‑level interpretability while reducing computational cost and improving generalization across heterogeneous datasets.
+Patient‑Level Evaluation
+Evaluation is performed at the patient level to reflect clinical use.
+For each patient $p_{m}$, window‑level probabilities $\hat{y}_{mi}$ from spectrograms $\mathcal{W}_{m}$ are aggregated by averaging,
 $\hat{y}_{m}=\frac{1}{|\mathcal{W}_{m}|}\sum_{i=1}^{|\mathcal{W}_{m}|}\hat{y}_{mi},$ (5)
 yielding a single prediction per patient.
-This procedure ensures that performance metrics correspond to generalization across distinct individuals rather than across overlapping signal segments, thereby aligning the evaluation with the outer cross‑validation design and providing clinically meaningful estimates of model utility.
+This ensures performance reflects generalization across individuals rather than overlapping signal segments, aligning evaluation with the outer cross‑validation design.
 IV Experiments and Results
-We evaluated our framework through a series of controlled experiments spanning multiple datasets, metrics, and model configurations.
-The following subsections detail the datasets used, evaluation criteria, baseline comparisons, and implementation choices, followed by analyses of model performance (Table I), aggregation strategies (Table II), and interpretability via Grad-CAM visualizations (Figure 3).
-IV-A Datasets
-We combined three independent EEG dataset’s resting state eyes closed subsets to assess robustness and generalizability: (i) Iowa, 64–channel BioSemi ActiveTwo recordings; (ii) University of New Mexico (UNM), multi-session 64–channel EEG from PD patients and controls; and (iii) San Diego, clinical 32–channel EEG representing hardware variability.
-All signals were aligned to our canonical montage set forth in III-A, with missing channels zero-filled to preserve indexing.
-This harmonization enabled cross-dataset evaluation without architectural modification.
-IV-B Evaluation Metrics
+Datasets
+We combined three independent public EEG datasets to assess robustness and cross‑dataset generalization: (i) Iowa [16], consisting of 64‑channel resting‑state recordings acquired under eyes‑open conditions; (ii) University of New Mexico (UNM) [16], comprising multi‑session 64‑channel EEG from Parkinson’s disease patients and controls with both eyes‑open and eyes‑closed resting‑state segments recorded within the same session; and (iii) San Diego [32], a clinical 32‑channel EEG dataset reflecting hardware and montage variability.
+For datasets providing medication annotations (UNM and San Diego), only recordings acquired in the on‑medication state were used.
+Evaluation Metrics
 With the dataset approximately balanced at the patient level, we report accuracy, area under the ROC curve (AUC), precision, and recall without adjustments.
-These metrics capture overall correctness, threshold discrimination, and the trade‑off between false positives and negatives, providing a robust evaluation under balanced conditions.
-IV-C Implementation and Reproducibility
-All experiments were implemented in Python with tensorflow/keras.
-Outputs were stored in DataFrames and logged to CSVs to ensure reproducibility.
-Training employed Adamax with an exponential learning‑rate schedule, early stopping, and checkpointing.
-Crucially, reproducibility is anchored by the nested cross‑validation channel selection routine (Algorithm 1), which formalizes how channels are chosen and evaluated across folds.
-Readers are referred to Algorithm 1 for the full specification.
+These metrics capture overall correctness, threshold discrimination, and the trade‑off between false positives and negatives, providing a robust evaluation of clinical relevance under balanced conditions for Parkinson’s detection.
+Implementation and Reproducibility
+All experiments were implemented in Python using tensorflow/keras, with results logged to CSVs for reproducibility.
+Training used Adamax with exponential decay, early stopping, and checkpointing.
+Reproducibility is ensured by the nested cross‑validation channel‑selection procedure (Algorithm 1), which governs principled channel selection evaluation across folds reducing the feature set; see Algorithm 1 and https://github.com/USD-AI-ResearchLab/Nested-Channel-Selection for the full specification and GitHub repository.
+Our Results
+Table I summarizes performance across three evaluation paradigms.
+The first, No Stratification, trains without patient‑level separation, permitting temporal and subject leakage; we also include results from Lee et al.
+[13] and Wu et al.
+[21], who used a comparable setup.
+The second, Single‑Population Blocking, reflects models optimized for a single site or cohort; our I‑4C model is shown alongside Anjum et al.
+[16] and Wu et al.
+[21] where they are evaluated with the same test set.
+On the other hand, Zhang et al.
+[29] tests on the UNM dataset and is included to reinforce single hold-out population performance.
+The third, Stratified Cross‑Validation, is our proposed framework, enforcing patient‑level separation across folds.
+Here, the All‑Channel model uses all electrodes, while reduced‑channel variants (1‑, 2‑, 4‑, 8‑, 16‑Channel) demonstrate principled channel selection.
+These comparisons reveal inflated performance under non‑stratified or singular population‑blocked designs and more clinically relevant estimates with multi-site stratified cross‑validation schema, ensuring realistic estimates.
 Algorithm 1 Nested Cross-Validation Channel Selection
@@
 27:return $\mathcal{C}^{*}$, $\bar{D}$
-IV-D Baselines
-Table I summarizes performance across three evaluation paradigms.
-The first, No Stratification, trains without patient‑level separation, permitting temporal and subject leakage; we also include results from Lee et al.
+IV-A Comparative Analysis
+The first paradigm, No Stratification, achieved the highest apparent performance (Acc = .985, AUC = .986, Prec = .995, Rec = .978), consistent with results reported by Lee et al.
 [13] and Wu et al.
-[21] , who used a comparable setup.
-The second, Single‑Population Blocking, reflects models optimized for a single site or cohort; our I‑4‑C model is shown alongside Anjum et al.
-[16] and Zhang et al.
-[29], both evaluated under similar conditions.
-The third, Stratified Cross‑Validation, is our proposed framework, enforcing patient‑level separation across folds.
-Here, the All‑Channel model uses all electrodes, while reduced‑channel variants (1‑, 2‑, 4‑, 8‑, 16‑Channel) demonstrate principled channel selection.
-These comparisons reveal inflated performance under non‑stratified or population‑blocked designs and more clinically relevant estimates with stratified cross‑validation.
-IV-E Comparative Analysis
-The first paradigm, No Stratification, achieved the highest apparent performance in our experiments (Acc = .985, AUC = .986, Prec = .995, Rec = .978), with Lee et al.
+[21].
+However, these scores reflect data leakage rather than genuine generalization, as patient overlap across folds inflates performance and near‑zero variance signals overfitting.
+We therefore include this paradigm only to illustrate how evaluation protocols that ignore temporal and patient‑level structure can yield misleading outcomes.
+Under Single‑Population Blocking, our 4‑Channel model achieved the strongest performance among evaluated methods.
+In the Iowa‑blocked evaluation, it attained Acc = .929, AUC = .929, Prec = .875, and Rec = 1.00, exceeding the mid‑.80s accuracies reported by Anjum et al.
 [16] and Wu et al.
-[21] reporting better metrics.
-However, these results reflect data leakage rather than genuine generalization since patient overlap across folds inflates performance, along with near‑zero variance that further signals overfitting.
-While the other architecture slightly outperforms ours under flawed evaluation, such comparisons are not meaningful: without patient‑level separation, performance cannot be interpreted as evidence of clinical utility.
-Under the Single‑Population Blocking paradigm, our 4‑Channel model decisively outperformed prior approaches.
-In the Iowa‑blocked evaluation, it achieved Acc = .929, AUC = .929, Prec = .875, and Rec = 1.00—substantially higher than the mid‑.80s accuracies reported by Anjum et al.
-[16], Zhang et al.
-[29], and Wu et al.
 [21] under comparable blocked designs.
-This margin underscores the strength of our feature extraction and channel selection strategy, even when models are tuned to a single dataset.
-However, such blocked evaluations can be misleading: by restricting training and testing to the same population, they risk inflating performance through site‑specific artifacts rather than genuine disease‑relevant features.
-Although population block testing offers a partial safeguard, it remains narrow, as models may still overfit to the unique characteristics of a single cohort.
-These limitations highlight why cross‑validation with patient‑level separation is essential for establishing true generalizability and clinical validity.
-Our framework is the first to provide clinically realistic estimates of generalization on heterogeneous EEG channel datasets by enforcing patient‑level separation through Stratified Cross‑Validation.
-Within this rigorous setting, the 4‑Channel configuration delivered the strongest overall trade‑off, achieving Acc = .806 $\pm$ .092 and Rec = .801 $\pm$ .134 while maintaining stable precision (.785 $\pm$ .159).
-The 2‑Channel model offered balanced but slightly weaker performance (Acc = .774, AUC = .797), whereas the 8‑Channel variant favored precision (.819 $\pm$ .108) at the expense of recall (.706 $\pm$ .157).
-By contrast, the 16‑Channel model underperformed across all metrics, suggesting that excessive electrodes introduce redundancy or noise.
-These results demonstrate not only the robustness of our channel selection strategy but also the importance of evaluating models under stratified, patient‑level cross‑validation, where inflated performance from site‑specific artifacts is minimized and true generalization can be assessed.
-Taken together, these results reveal that inflated scores often stem from temporal or patient leakage rather than true predictive ability, and that best‑case performance within a single dataset can mask poor generalization under heterogeneous conditions.
-By contrast, our stratified framework enforces patient‑level separation, yielding more conservative but clinically meaningful estimates.
-Moderate channel configurations consistently provided the most reliable balance, whereas both minimal and excessive channel counts degraded performance, underscoring the value of principled selection in the inner loop.
-Overall, findings highlight the need of stratification and the unique contribution of our framework in aligning evaluation with clinical deployment.
-IV-F Patient-Level Aggregation Ablation
-Table II presents an ablation of aggregation strategies for patient-level decisions in our nested cross-validation framework.
-Across all configurations, Mean and Median aggregation consistently produced the most balanced outcomes, reinforcing their reliability for patient‑level evaluation.
-The 4‑Channel models performed best overall, with both rules yielding strong accuracy and recall, while 2‑Channel models also improved on the baseline.
-Other strategies showed less stability: Majority performed competitively but leaned toward higher precision, Geometric Mean (GMean) was slightly weaker yet serviceable, and Max and Min proved unreliable—either overemphasizing precision at the cost of recall or collapsing under outliers.
-Importantly, the Iowa Blocked Test followed this same general trend: Mean, Median, and Majority aggregation produced the strongest and most stable results, while Min and Max underperformed.
-Overall, these findings show that aggregation is not a secondary detail but a central design choice that directly shapes clinical validity.
-Robust rules such as Mean and Median not only stabilize fold‑to‑fold variance but also preserve the sensitivity–specificity balance needed for screening contexts.
-Conversely, unstable rules amplify the same risks seen with poor stratification or excessive channel counts; namely, poor generalization.
-Thus, careful aggregation complements stratified cross‑validation and principled channel selection, forming our framework for reliable patient‑level evaluation.
-TABLE I: Performance comparison under different stratification strategies and channel subsets.
+This suggests effective feature extraction and channel selection under single‑cohort tuning; however, restricting both training and testing to a single population remains vulnerable to site‑specific or acquisition‑dependent artifacts and does not ensure generalization beyond the evaluated cohort [29].
+Accordingly, strong performance under population‑blocked testing should be interpreted cautiously.
+In contrast, our framework provides clinically realistic estimates of generalization on heterogeneous EEG datasets by enforcing patient‑level separation through Stratified Cross‑Validation.
+Within this setting, the 4‑Channel configuration achieved the strongest overall trade‑off (Acc = .806 $\pm$ .092, Rec = .801 $\pm$ .134) while maintaining stable precision (.785 $\pm$ .159).
+The 2‑Channel model showed balanced but weaker performance, whereas the 8‑Channel variant favored precision at the expense of recall; the 16‑Channel model underperformed across all metrics, reflecting a trade‑off between spatial coverage and channel reliability under patient‑level evaluation.
+Importantly, channel selection is driven by stability‑adjusted inner‑fold performance: as the retained subset grows, marginal contributions diminish and variability across folds increases, leading to reduced generalization under patient‑level cross‑validation.
+Taken together, these results show that inflated scores often arise from temporal or patient leakage, and that best‑case performance within a single dataset can obscure poor generalization under heterogeneous conditions.
+By contrast, stratified evaluation with principled channel selection yields more conservative but clinically meaningful estimates aligned with deployment‑ready generalization.
+IV-B Patient‑Level Aggregation Ablation
+Table II presents an ablation of patient‑level aggregation strategies within our nested cross‑validation framework.
+Across configurations, Mean and Median aggregation consistently produced the most balanced and stable outcomes, particularly for the 4‑Channel model, which achieved the strongest overall accuracy and recall.
+The 2‑Channel models showed similar aggregation behavior, indicating that aggregation strategies generalize across folds: Majority performed competitively but favored precision, GMean was modestly weaker, and Max and Min proved unstable, either amplifying outliers or collapsing recall.
+The Iowa Blocked Test followed the same pattern, with Mean, Median, and Majority yielding the most consistent results across folds.
+Overall, these findings indicate that aggregation is a central design choice shaping clinical validity rather than a secondary detail.
+Robust rules such as Mean and Median stabilize fold‑to‑fold variance while preserving the sensitivity–specificity balance required for screening contexts.
+In contrast, unstable aggregation strategies exacerbate the same generalization failures observed under poor stratification or excessive channel counts.
+Accordingly, careful aggregation complements stratified cross‑validation and principled channel selection, completing our framework for reliable, clinically relevant, patient‑level evaluation.
+TABLE I: Performance under different stratification and channel configs.
 Model Acc $\sigma$ AUC $\sigma$ Prec $\sigma$ Rec $\sigma$
@@
 16-Channel .748 .092 .753 .143 .793 .112 .703 .154
-TABLE II: Ablation of aggregation strategies for patient-level decision making under different channel configurations.
+TABLE II: Ablation of aggregates for patient-level decisions.
 Method Acc $\sigma$ AUC $\sigma$ Prec $\sigma$ Rec $\sigma$
@@
 The model emphasizes recurrent theta activity (4–8 Hz), with secondary attention in alpha (8–12 Hz) and occasional beta (12–30 Hz).
-These theta–alpha patterns broaden and narrow rhythmically, producing a pulsing appearance distinct from noise or artifacts.
-While edge effects (e.g., aliasing, zero‑padding) sometimes attract attention, the dominant focus remains on physiologically plausible oscillations, underscoring both the strengths and limits of convolutional interpretability
-IV-G Grad-CAM Interpretability Visualization
-Grad-CAM overlays in Figure 3 revealed that the network consistently emphasized the theta band (4–8 Hz) across channels, with secondary contributions from alpha (8–12 Hz) and beta (12–30 Hz) ranges.
+These theta–alpha patterns broaden and narrow, producing a pulsing appearance distinct from noise or artifacts.
+While edge effects (e.g., aliasing, zero‑padding) sometimes attract attention, the dominant focus remains on physiologically plausible oscillations.
+IV-C Grad-CAM Interpretability Visualization
+Grad-CAM overlays in Figure 3 reveal that the network consistently emphasized the theta band (4–8 Hz) across channels, with secondary contributions from alpha (8–12 Hz) and beta (12–30 Hz) ranges.
 In several cases, the model highlighted recurrent theta–alpha structures with rhythmic fluctuations in spectral bandwidth, producing a pulsing appearance distinct from artifacts.
@@
 The frequency-specific spatially distributed activations across channels further indicate that the network captured cross-channel dependencies rather than isolated channel effects.
-This pattern aligns with prior EEG studies in PD, which report consistent alterations in theta and beta.
+This pattern aligns with prior EEG studies in PD, which report consistent alterations in theta and beta bands.
 For example, increased frontal theta activity and impaired beta desynchronization during lower-limb movement have been linked to motor deficits and freezing of gait (FOG) in PD patients [33].
@@
 The inner loop enabled principled channel selection, reducing dimensionality and improving generalization by emphasizing the most informative electrodes.
-Patient‑level aggregation, particularly mean and median, further reinforced clinical reliability.
+Patient‑level aggregation, particularly mean, median, and majority, further reinforced clinical reliability.
 Collectively, these components show nested cross-validation as a principled model‑agnostic solution to leakage, dimensionality, and heterogeneity challenges that have historically inflated performance.
-Beyond PD, our architecture offers a reproducible framework for EEG and ML studies emphasizing scientific rigor, transparency, and clinical applicability.
+Beyond PD, our architecture offers a reproducible framework for EEG and ML studies emphasizing scientific rigor and clinical applicability.
 Acknowledgments
-The author acknowledges Microsoft Copilot for assistance with grammar and structural edits, contextual extrapolation in the results section, and figure preparation with extensive human input [32].
+This work was supported by the National Science Foundation under Grant No.
+#2346643, the U.S.
+Department of Defense under Award No.
+#FA9550-23-1-0495, and the U.S.
+Department of Education under Grant No.
+P116Z240151.
+Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, the U.S.
+Department of Defense, or the U.S.
+Department of Education.
+The author acknowledges Microsoft Copilot for assistance with grammar, structural edits, contextual extrapolation, and figure preparation, all with extensive human input [34].
 All core ideas, analyses, and interpretations are the author’s own based on previous literature and experimental analysis, as cited in the manuscript.